7. Show that σ2 E(X-0 and Var(X if X1, . . . , Xn are independent and identically distributed with E(Xi) = 0 and E(X2) = σ2 for i = 1,-.. , n
Answer #1

Proof 1:

We know,

\bar X = \frac{\sum_{i = 1}^{n}X_i}{n}

Now,

E(\bar X) = E(\frac{\sum_{i = 1}^{n}X_i}{n}) = \frac{\sum_{i = 1}^{n}E(X_i)}{n} = 0 As given, E(Xi) = 0

Proof 2:

Var(X_i) = E(X_i - E(X_i))^2 = E(X_i^2) - [E(X_i)]^2 = E(X_i^2) = \sigma^2, As E(Xi) = 0

We know,

\bar X = \frac{\sum_{i = 1}^{n}X_i}{n}

Var(\bar X) = Var(\frac{\sum_{i = 1}^{n}X_i}{n}) = \frac{\sum_{i = 1}^{n}Var(X_i)}{n^2} = \frac{\sum_{i = 1}^{n}\sigma^2}{n^2} = \frac{n \sigma^2}{n^2} = \frac{\sigma^2}{n} (Proved)

** If any confusion, please comment.